### v2

# An Objective Improvement Approach

Abstract.
While discounted payoff games and classic games that reduce to them, like parity and mean-payoff games, are symmetric, their solutions are not. We have taken a fresh view on the properties that optimal solutions need to have, and devised a novel way to converge to them, which is entirely symmetric. We achieve this by building a constraint system that uses every edge to define an inequation, and update the objective function by taking a single outgoing edge for each vertex into account. These edges loosely represent strategies of both players, where the objective function intuitively asks to make the inequation to these edges sharp. In fact, where they are not sharp, there is an ‘error’ represented by the difference between the two sides of the inequation, which is 0 where the inequation is sharp. Hence, the objective is to minimise the sum of these errors. For co-optimal strategies, and only for them, it can be achieved that all selected inequations are sharp or, equivalently, that the sum of these errors is zero. While no co-optimal strategies have been found, we step-wise improve the error by improving the solution for a given objective function or by improving the objective function for a given solution. This also challenges the gospel that methods for solving payoff games are either based on strategy improvement or on value iteration.
Key words and phrases:
Discounted Payoff Games, Payoff Games, Objective Improvement, Strategy Improvement
1. Introduction
We study turn-based zero sum games played between two players on directed graphs. The two players take turns to move a token along the vertices of a finite labelled graph with the goal to optimise their adversarial objectives.
Various classes of graph games are characterised by the objective of the players, for instance in parity games the objective is to optimise the parity of the dominating colour occurring infinitely often, while in discounted and mean-payoff games the objective of the players is to minimise resp. maximise the discounted and limit-average sum of the colours.
Solving graph games is the central and most expensive step in many model checking [Koz83, EJS93, Wil01, dAHM01, AHK02, SF06a], satisfiability checking [Koz83, Var98, Wil01], and synthesis [Pit06, SF06b] algorithms. Progress in algorithms for solving graph games will therefore allow for the development of more efficient model checkers and contribute to bringing synthesis techniques to practice.
There is a hierarchy among the graph games mentioned earlier, with simple and well known reductions from parity games to mean payoff games, from mean-payoff games to discounted payoff games, and from discounted payoff games to simple stochastic games like the ones from [ZP96], while no reductions are known in the other direction. Therefore, one can solve instances of all these games by using an algorithm for stochastic games. All of these games are in UP and co-UP [Jur98], while no tractable algorithm is known.
Most research has focused on parity games: as the most special class of games, algorithms have the option to use the special structure of their problems, and they are most directly linked to the synthesis and verification problems mentioned earlier. Parity games have thus enjoyed a special status among graph games and the quest for efficient algorithms [EL86, EJ91, McN93, ZP96, BCJ+97, Zie98, Obd03, BDM16, BDM18] to solve them has been an active field of research during the last decades, which has received further boost with the arrival of quasi-polynomial techniques [JL17, FJdK+19, LB20, LPSW20, DS22, CJK+22, BDM+24].
Interestingly, the one class of efficient techniques for solving parity games that does not (yet) have a quasi-polynomial approach is strategy improvement algorithms [Lud95, Pur95, VJ00, BV07, Sch08, Fea10, STV15], a class of algorithms closely related to the Simplex for linear programming, known to perform well in practice. Most of these algorithms reduce to mean [BV07, Sch08, STV15, BDM20] or discounted [Lud95, Pur95, FGO20, Koz21] payoff games.
With the exception of the case in which the fixed-point of discounted payoff games is explicitly computed [ZP96], all these algorithms share a disappointing feature: they are inherently non-symmetric approaches for solving an inherently symmetric problem. However, some of these approaches have a degree of symmetry. Recursive approaches treat even and odd colours symmetrically, one at a time, but they treat the two players very differently for a given colour. Symmetric strategy improvement [STV15] runs a strategy improvement algorithms for both players in parallel, using the intermediate results of each of them to inform the updates of the other, but at heart, these are still two intertwined strategy improvement algorithms that, individually, are not symmetric. This is in due to the fact that applying strategy improvement itself symmetrically can lead to cycles [Con93].
The key contribution of this paper is to devise a new class of algorithms to solve discounted payoff games, which is entirely symmetric. Like strategy improvement algorithms, it seeks to find co-optimal strategies, and improves strategies while they are not optimal. In order to do so, however, it does not distinguish between the strategies of the two players. This seems surprising, as maximising and minimising appear to pull in opposing directions.
Similar to strategy improvement approaches, the new objective improvement approach turns the edges of a game into constraints (here called inequations), and minimises an objective function. However, while strategy improvement algorithms take only the edges in the strategy of one player (and all edges of the other player) into account and then finds the optimal response by solving the resulting one player game, objective improvement always takes all edges into account. The strategies under consideration then form a subset of the inequations, and the goal would be to make them sharp (i.e. strict, tight, satisfied as equations), which only works when both strategies are optimal. When they are not, then there is some offset for each of the inequations, and the objective is to reduce this offset in every improvement step. This treats the strategies of both players completely symmetrically.
Organisation of the Paper.
The paper is organised as follows. After the preliminaries in Section 2, we start by outlining our method and use a simple game to explain it in Section 3. We then formally introduce our objective improvement algorithm in Section 4, keeping the question of how to choose better strategies abstract. Section 5 discusses how to find better strategies, and Section 6 provides an experimental evaluation of the algorithm. We finally wrap up with a discussion of our results in Section 7.
2. Preliminaries
A discounted payoff game (DPG) is a tuple $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$, where $V=V_{\min}\cup V_{\max}$ are the vertices of the game, partitioned into two disjoint sets $V_{\min}$ and $V_{\max}$, such that the pair $(V,E)$ is a finite directed graph without sinks. The vertices in $V_{\max}$ (resp, $V_{\min}$) are controlled by Player Max or maximiser (resp, Player Min or minimiser) and $E\subseteq V\times V$ is the edge relation. Every edge has a weight represented by the function $w:E\to\mathbb{R}$, and a discount factor represented by the function $\lambda:E\to[0,1)$. When the discount factor is uniform, i.e. the same for every edge, it is represented by a constant value $\lambda\in[0,1)$. For ease of notation, we write $w_{e}$ and $\lambda_{e}$ instead of $w(e)$ and $\lambda(e)$. A play on $\mathcal{G}$ from a vertex $v$ is an infinite path, which can be represented as a sequence of edges $\rho=e_{0}e_{1}e_{2}\ldots$ such that, for every $i\in\mathbb{N}$, $e_{i}=(v_{i},v_{i+1})\in E$ and $v_{0}=v$. By $\rho_{i}$ we refer to the $i$-th edge of the play. The outcome of a discounted game $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ for a play $\rho$ is $\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\prod_{j=0}^{i-1}\lambda_{e_{j}}$. For games with a constant discount factor, this simplifies into $\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\lambda^{i}$.
A positional strategy for Max is a function $\sigma_{\max}:V_{\max}\to V$ that maps each Max vertex to a vertex according to the set of edges, i.e. $(v,\sigma_{\max}(v))\in E$. Positional Min strategies are defined accordingly, and we call the set of positional Min and Max strategies $\Sigma_{\min}$ and $\Sigma_{\max}$, respectively.
A pair of strategies $\sigma_{\min}$ and $\sigma_{\max}$, one for each player, defines a unique play $\rho(v,\sigma_{\min},\sigma_{\max})$ from each vertex $v\in V$. Discounted payoff games are positionally determined [ZP96]:
$\adjustlimits{\max}_{\sigma_{\max}\in\Sigma_{\max}}{\min}_{\sigma_{\min}\in\Sigma_{\min}}\mathsf{out}(\rho(v,\sigma_{\min},\sigma_{\max}))=\adjustlimits{\min}_{\sigma_{\min}\in\Sigma_{\min}}{\max}_{\sigma_{\max}\in\Sigma_{\max}}\mathsf{out}(\rho(v,\sigma_{\min},\sigma_{\max}))$
holds for all $v\in V$, and neither the strategy, nor the value computed from the strategy, changes when we allow more powerful classes of strategies that allow for using memory and/or randomisation for one or both players.
The value of a minimiser strategy $\sigma$, denoted by $\mathsf{val}_{\sigma}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{val}_{\sigma}:v\mapsto\max_{\sigma_{\max}\in\Sigma_{\max}}\mathsf{out}(\rho(v,\sigma,\sigma_{\max}))\ ,$
The value of a maximiser strategy $\sigma$, denoted by $\mathsf{val}_{\sigma}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{val}_{\sigma}:v\mapsto\min_{\sigma_{\min}\in\Sigma_{\min}}\mathsf{out}(\rho(v,\sigma_{\min},\sigma))\ .$
The resulting value of $\mathcal{G}$, denoted by $\mathsf{val}_{\mathcal{G}}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{val}_{\mathcal{G}}:v\mapsto\max_{\sigma_{\max}\in\Sigma_{\max}}\min_{\sigma_{\min}\in\Sigma_{\min}}\mathsf{out}(\rho(v,\sigma_{\min},\sigma_{\max}))\ .$
A positional maximiser (resp. minimiser) strategy $\sigma$ is said to be optimal if, and only if, $\mathsf{val}_{\mathcal{G}}(v)=w_{(v,\sigma(v))}+\lambda_{(v,\sigma(v))}\mathsf{val}_{\mathcal{G}}(\sigma(v))$ holds for all maximiser (resp. minimiser) vertices.
Likewise, we define the value of a pair of strategies $\sigma_{\min}$ and $\sigma_{\max}$, denoted $\mathsf{val}_{\sigma_{\min},\sigma_{\max}}:V\rightarrow\mathbb{R}$, as $\mathsf{val}_{\sigma_{\min},\sigma_{\max}}:v\mapsto\mathsf{out}(\rho(v,\sigma_{\min},\sigma_{\max}))$.
As we treat both players symmetrically in this paper, we define a pair of strategies $\sigma:V\mapsto V$ whose restriction to $V_{\min}$ and $V_{\max}$ are a minimiser strategy $\sigma_{\min}$ and a maximiser strategy $\sigma_{\max}$, respectively. We then write $\rho(v,\sigma)$ instead of $\rho(v,\sigma_{\min},\sigma_{\max})$ and $\mathsf{val}_{\sigma}$ instead of $\mathsf{val}_{\sigma_{\min},\sigma_{\max}}$. To ease the reading, when the strategy is clear from the context, we shall drop the subscript.
If both of these strategies are optimal, we call $\sigma$ a joint co-optimal strategy. This is the case if, and only if, $\mathsf{val}_{\mathcal{G}}=\mathsf{val}_{\sigma}$ holds.
We define a solution to a discounted payoff game $\mathcal{G}$ to be a valuation $\mathsf{val}$ for the vertices such that, for every edge $e=(v,v^{\prime})$, it holds that [^1] These are the constraints represented in $H$ in Section 4.
  • $\mathsf{val}(v)\leq w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ if $v$ is a minimiser vertex and
  • $\mathsf{val}(v)\geq w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ if $v$ is a maximiser vertex.
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution.
Note that we are interested in the value of each vertex, not merely if the value is greater or equal than a given threshold value.
2.1. Simplex method
A common way to compute $\mathsf{val}_{\mathcal{G}}$ is to create a linear programming instance from $\mathcal{G}$. In this paragraph we provide the background for solving a linear programming instance and outline the Simplex method that is the standard algorithm used to find a solution.
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that maximises (or minimises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Every linear programming instance can be expressed in a simpler standard form by applying transformations to the objective function and to the constraints. A standard form requires to find $x$ that minimises $c\cdot x$ such that $A\cdot x=b$ and $x\geq 0$, with $b$ positive. Given a game $\mathcal{G}$, we can define an objective function, the input matrix $A$, and the vectors $b$ and $c$, such that $x=\mathsf{val}_{\mathcal{G}}$, in the following way. $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a non-zero value (with the exception of self-loops that have only one non-zero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}+a_{ik}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$, $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative for the minimiser, and the objective function maximises the sum of the values of $x$.
From a graphical point of view, the set of costraints, i.e. the inequations represented by $A$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope is a solution. The objective function, however, provides an additional vector that shows the direction of better solutions. As a result, if there is an optimal solution (and there always is an optimal solution in the linear programming problems we consider in this paper), then there is an optimal solution that lies on a corner of the polytope. This observation allows to restrict the search space for optimal solutions to the corners of the polytope. In fact, it suffices to move along the edges of the polytope to find and optimal solution.
From an algebraic point of view, each vector $x$ that corresponds to a corner of the polytope satisfies $n$ constraints (vectors) linearly independent of the linear programming problem. Such $x$ is called basis since every constraint of the problem can be written as a unique combination of the $n$ constraints satisfied by $x$. The associated set of constraints to the basis is called $B$. Hence, to find the optimal solution of the linear programming problem one has to find the basis (admissible solution) that minimises the objective function (optimality).
The Simplex method assumes that the problem is in the standard form and solves it by finding a basis and then by moving to an adjacent basis that improves the value of the objective function. Two basis are adjacent if in the graphical representation they correspond to adjacent corners. This implies that the set of constraints they satisfy differ for one single vector. The algorithm stops when the adjacent bases do not improve the value of the objective function of the basis currently selected.
input : The elements of a linear programming problem $A,b,c$
output : The optimal solution vector $x$
1 $B\leftarrow{}$ random initial basis matrix
2 while true do
    foreach $j$ non basis index do
       3 $c^{\prime}_{j}\leftarrow{}$ reduced cost over objective function
    end foreach
   4if $c^{\prime}>0$ then
       return $x$
    end if
   5Select $j$ such that $c^{\prime}_{j}<0$
   6 $B\leftarrow{}\mathsf{Pivot}(A,b,c,j)$
   7 $x\leftarrow{}\mathsf{Update}(B,c)$
end while
Algorithm 1 Simplex method
The procedure is reported in Algorithm 1. At Line 1, an initialisation step selects a first basis. The “for each” loop at Line 3 computes the difference of the current objective function value and the one obtained by selecting a different constraint $j$. In case all these differences are positive, then the current solution cannot be improved and it is returned at Line 4. Otherwise, one of the improving adjacent corner is selected (Line 5). The move from one basis to an adjacent one is called pivoting operation (Line 6). Once the basis matrix has been changed, the solution vertex can be updated (Line 7).
3. Outline and Motivation Example
We start by considering the simple discounted payoff game of Figure 1, assuming that it has some uniform discount factor $\lambda\in[0,1)$. In this game, the minimiser (who owns the right vertex, marked by a square), has only one option: she always has to use the self-loop, which earns her an immediate reward of $1$. The overall reward the minimiser reaps for a play that starts in her vertex is therefore $1+\lambda+\lambda^{2}+\ldots=\frac{1}{1-\lambda}$.
The maximiser (who owns the left vertex, $b$, marked by a circle) can choose to either use the self-loop, or to move to the minimiser vertex (marked by a square), both yielding no immediate reward.
ba00001111
Figure 1. A discounted Payoff Game. Maximiser vertices are depicted as a circle, minimizer ones as a square.
If the maximiser decides to stay forever in his vertex (using the self-loop), his overall reward in the play that starts at (and, due to his choice, always stays in) his vertex, is 00. If he decides to move on to the minimiser vertex the $n^{th}$-time, then the reward is $\frac{\lambda^{n}}{1-\lambda}$.
The optimal decision of the maximiser is therefore to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}(v)=w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ holds [ZP96]; these edges correspond to the optimal decisions for the respective player.
For our running example game of Figure 1 with a fixed discount factor $\lambda\in[0,1)$, the constraints a solution must satisfy are:
  1. (1)
    $\mathsf{val}(a)\leq 1+\lambda\mathsf{val}(a)$   for the self-loop of the minimiser vertex;
  2. (2)
    $\mathsf{val}(b)\geq\lambda\mathsf{val}(b)$         for the self-loop of the maximiser vertex; and
  3. (3)
    $\mathsf{val}(b)\geq\lambda\mathsf{val}(a)$        for the transition from the maximiser to the minimiser vertex.
The unique valuation that satisfies these inequations and produces a sharp inequation (i.e. satisfied as equation) for some outgoing edge of each vertex assigns $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=\frac{\lambda}{1-\lambda}$. This valuation also defines the optimal strategies of the players (to stay for the minimiser, and to move on for the maximiser).
Solving a discounted payoff game means finding this valuation and/or these strategies.
We discuss a symmetric approach to find this unique valuation. Our approach adjusts linear programming in a natural way that treats both players symmetrically: we maintain the set of inequations for the complete time, while approximating the goal of “one equality per vertex” by the objective function. To do that, we initially fix an arbitrary outgoing edge for every vertex (a strategy), and minimise the sum of the distances between the left and right side of the inequations defined by these edges, which we call the offset of this edge. This means, for an edge $e=(v,v^{\prime})$, to minimise the difference between $\mathsf{val}(v)$ (left side of the inequation) and $w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ (right side).
To make this clear, we consider again the example of Figure 1 and use both self loops as the strategies for the players fixed at the beginning in our running example. The offset for the selected outgoing edge of the minimiser vertex $a$ is equal to $1-(1-\lambda)\mathsf{val}(a)$, while the offset for the selected outgoing edge of the maximiser vertex $b$ is equal to $(1-\lambda)\mathsf{val}(b)$. The resulting overall objective consists, therefore, in minimising the value $1-(1-\lambda)\mathsf{val}(a)+(1-\lambda)\mathsf{val}(b)$.
This term is always non-negative, since it corresponds to the sum of the edges contributions that are all non-negative. Moreover, when only optimal strategies are selected to form this objective function, the value 00 can be taken, and where it is taken, it defines the correct valuation.
As the maximiser’s choice to take the self-loop is not optimal, the resulting objective function the strategies define, that is $1-(1-\lambda)\mathsf{val}(a)+(1-\lambda)\mathsf{val}(b)$, cannot reach 00. But let us take a look at what an optimal solution w.r.t. this objective function looks like.
Optimal solutions can be taken from the corners of the polytope defined by the inequations, also known as its simplex. In this case, the optimal solution (w.r.t. this initial objective function) is defined by making inequations (1) and (3) sharp: this provides the values $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=\frac{\lambda}{1-\lambda}$; the objective function takes the value $\lambda$ at this point.
For comparison, in the other corner of the simplex, defined by making inequations (2) and (3) sharp, we obtain the values $\mathsf{val}(a)=\mathsf{val}(b)=0$; the objective function takes the value $1$ at this point. Finally, if we consider the last combination, making (1) and (2) sharp, this provides the values $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=0$, so that inequation (3) is not satisfied; this is therefore not a corner of the simplex.
Thus, in this toy example, while selecting the wrong edge cannot result in the objective function taking the value 00, we still found the optimal solution. In general, we might need to update the objective function. To update the objective function, we change the outgoing edges of some (or all) vertices, such that the overall value of the objective function goes down. Note that this can be done not only when the linear program returns an optimal solution, but also during its computation. For example, when using a simplex method, updating the objective function can be used as an alternative pivoting rule at any point of the traversal of the simplex.
In general, the valuation returned as solution is computed by using objective functions based on sub-optimal strategies. The simplest way of seeing non-optimal strategies is to use different discount factors [^2] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1, let’s say $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three equations are: (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$, (2) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(b)$, and (3) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(a)$. Making the adjusted inequations (2) and (3) sharp still results in the values $\mathsf{val}(a)=\mathsf{val}(b)=0$, and the objective function still takes the value of $1$. While making inequations (1) and (3) sharp provides the values $\mathsf{val}(a)=3$ and $\mathsf{val}(b)=2$; the objective function takes the value $\frac{4}{3}$ at this point. Finally, if we consider the last combination, making (1) and (2) sharp still conflicts with inequation (3).
Thus, $\mathsf{val}(a)=\mathsf{val}(b)=0$ would be the optimal solution for the given objective function, which is not the valuation of the game. We will then update the candidate strategies so that the sum of the offsets goes down.
3.1. Comparison with strategy improvement
The closest relatives to our new approach are strategy improvement algorithms. Classic strategy improvement approaches solve the problem of finding the valuation of a game (and usually also co-optimal strategies) by (1) fixing a strategy for one of the players (we assume w.l.o.g. that this is the maximiser), (2) finding a valuation function for the one player game that results from fixing this strategy (often together with an optimal counter strategy for their opponent), and (3) updating the strategy of the maximiser by applying local improvements. This is repeated until no local improvements are available, which entails that the constraint system is satisfied.
input : A discounted payoff game $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$
output : The valuation $\mathsf{val}$ of $\mathcal{G}$
1 $f\leftarrow{}\mathsf{ObjectiveFunction}(\mathcal{G})$
2 $\sigma\leftarrow{}\mathsf{ChooseInitialStrategy}(\mathcal{G})$
3 while true do
   4 $H\leftarrow{}\mathsf{Inequations}(\mathcal{G},\sigma)$
   5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f)$
   6 if $\sigma$ is optimal then
       return $\mathsf{val}$
    end if
[…]
end while
Algorithm 2 Strategy Improvement
These steps can be identified in Algorithm 2. The objective function is fixed at the beginning (Line 1). It simply maximises the sum of the values of all the vertices, and thus implicitly the value of all vertices. (1) The strategy of one of the two player (for convention the maximiser – if we want to instead fix the strategy of the minimiser, we would have to minimise the sum of the values for all vertices instead) is initialised at Line 2. (2) Before the computation of the valuation at Line 5, the set of constraints is updated accordingly to the selected strategy of the maximiser (Line 4). At this point, if the strategy is optimal w.r.t. the current valuation, then the algorithm ends. Otherwise, (3) there exist an improvement, which is applied at Line 7.
For Step (2) of this approach, we can use linear programming, which does invite a comparison with our technique. The linear program for solving Step (2) would not use all inequations: it would, instead, replace the inequations defined by the currently selected edges of the maximiser by equations, while dropping the inequations defined by the other maximiser transitions. The objective function would then be to maximise the values of all vertices while still complying with the remaining (in)equations.
Thus, in our novel symmetric approach the constraints remain while the objective is updated, in strategy improvement, instead, the objective remains while the constraints are updated. Moreover, the players and their strategies are treated quite differently in strategy improvement algorithms: while the candidate strategy of the maximiser results in making the inequations of the selected edges sharp (and dropping all other inequations of maximiser edges), the optimal counter strategy is found by minimising the objective. This is again in contrast to our novel symmetric approach, which treats both players equally.
A small further difference is in the valuations that can be taken: the valuations that strategy improvement algorithms can take are the valuations of strategies, while the valuations our objective improvement algorithm can take on the way are the corners of the simplex defined by the inequations. Except for the only intersection point between the two (the valuation of the game), these corners of the simplex do not relate to the value of strategies. Table 1 summarises these observations.
Objective Improvement Strategy Improvement
players symmetric treatment asymmetric treatment
constraints remain the same: change:
one inequation per edge one inequation for each edge
defined by the current strategy for
the strategy player, one inequation
for every edge of their opponent
objective minimise errors for selected edges maximise values
update objective: strategy:
one edge for each vertex one edge for each vertex
of the strategy player
valuations corners of simplex defined by strategies
Table 1. A comparison of the novel objective improvement with classic strategy improvement.
4. General Objective Improvement
In this section, we present the approach outlined in the previous section more formally, while keeping the most complex step – updating the candidate strategy to one which is better in that it defines an optimisation function that can take a smaller value – abstract. (We turn back to the question of how to find better strategies in Section 5.) This allows for discussing the principal properties more clearly.
input : A discounted payoff game $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$
output : The valuation $\mathsf{val}$ of $\mathcal{G}$
1 $H\leftarrow{}\mathsf{Inequations}(\mathcal{G})$
2 $\sigma\leftarrow{}\mathsf{ChooseInitialStrategies}(\mathcal{G})$
3 while true do
   4 $f_{\sigma}\leftarrow{}\mathsf{ObjectiveFunction}(\mathcal{G},\sigma)$
   5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
   6 if  $f_{\sigma}(\mathsf{val})=0$ then
       return $\mathsf{val}$
    end if
[…]
end while
Algorithm 3 Objective Improvement
A general outline of our objective improvement approach is reported in Algorithm 3. Before describing the procedures called by the algorithm, we first outline the principle.
When running on a discounted payoff game $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$, the algorithm uses a set of inequations defined by the edges of the game and the owner of the source of each edge. This set of inequations denoted by $H$ contains one inequation for each edge and (different to strategy improvement approaches whose set of inequations is a subset of $H$) $H$ never changes.
The inequations from $H$ are computed by a function called $\mathsf{Inequations}$ (Line 1) that, given the discounted game $\mathcal{G}$, returns the set made of one inequation per edge $e=(v,v^{\prime})\in E$, defined as follows:
$I_{e}=\begin{cases}\mathsf{val}(v)\geq w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})&\text{if }v\in V_{\max},\\ \mathsf{val}(v)\leq w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})&\text{otherwise.}\end{cases}$
The set $H=\{I_{e}\mid e\in E\}$ is defined as the set of all inequations for the edges of the game.
The algorithm also handles strategies for both players, treated as a single strategy $\sigma$. They are initialised (for example randomly) by the function $\mathsf{ChooseInitialStrategies}$ at Line 2.
This joint strategy is used to define an objective function $f_{\sigma}$ by calling function $\mathsf{ObjectiveFunction}$ at Line 4, whose value on an evaluation $\mathsf{val}$ is: $f_{\sigma}(\mathsf{val})=\sum_{v\in V}f_{\sigma}(\mathsf{val},v)$ with the following objective function components:
$f_{\sigma}(\mathsf{val},v)=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$
where the offset of an edge $(v,v^{\prime})$ for a valuation is defined as follows:
$\mathsf{offset}(\mathsf{val},(v,v^{\prime}))=\begin{cases}\mathsf{val}(v)-(w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}(v^{\prime}))&\text{if }v\in V_{\max},\\ (w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}(v^{\prime}))-\mathsf{val}(v)&\text{otherwise.}\end{cases}$
This objective function $f_{\sigma}$ is given to a linear programming algorithm, alongside with the inequations set $H$. We underline that, due to the inequation $I_{(v,v^{\prime})}$, the value of $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))$ is non-negative for all $(v,v^{\prime})\in E$ in any valuation $\mathsf{val}$ (optimal or not) that satisfies the system of inequations $H$.
{obs}
At Line 6 of Algorithm 3, the value of $f_{\sigma}(\mathsf{val})$ is non-negative.
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, accordingly to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We denote with $\mathsf{val}_{\mathbf{b}}$ the evaluation of $\mathbf{b}$. As a consequence, the call $\mathsf{LinearProgramming}(H,f_{\sigma})$ to some linear programming algorithm returns a valuation $\mathsf{val}$ of the vertices that minimise $f_{\sigma}$ while satisfying $H$ (Line 5), and for convenience require this valuation to also be $\mathsf{val}_{\mathbf{b}}$ for some basis $\mathbf{b}$ of $H$. (Note that the simplex algorithm, for example, only uses valuations of this form in every step.) We call this valuation a valuation associated to $\sigma$.
We say that a valuation $\mathsf{val}$ defines strategies of both players if, for every vertex $v\in V$, the inequation of (at least) one of the outgoing edges of $v$ is sharp. These are the strategies defined by using, for every vertex $v\in V$, an outgoing edge for which the inequation is sharp. Note that there can be more than one of these inequations for some of the vertices.
{obs}
If, for a solution $\mathsf{val}$ of $H$, $f_{\sigma}(\mathsf{val})=0$ holds, then, for every vertex $v\in V$, the inequation $I_{(v,\sigma(v))}$ for the edge $(v,\sigma(v))$ is sharp, and $\mathsf{val}$ therefore defines strategies for both players – those defined by $\sigma$, for example.
Instead of checking whether $\mathsf{val}$ defines strategies of both players, we can use, alternatively, $f_{\sigma}(\mathsf{val})=0$ as a termination condition, as shown at Line 6 in Algorithm 3, since in this case $\sigma$ must define co-optimal strategies.
Theorem 1.
If $\sigma$ describes co-optimal strategies, then $f_{\sigma}(\mathsf{val})=0$ holds at Line 6 of Algorithm 3. If $\mathsf{val}$ (from Line 5 of Algorithm 3) defines joint strategies $\sigma^{\prime}$ for both players, then $\sigma^{\prime}$ is co-optimal and $\mathsf{val}$ is the valuation of $\mathcal{G}$.
Proof 4.1.
The valuation $\mathsf{val}=\mathsf{val}_{\mathcal{G}}$ of the game is the unique solution of $H$ for which, for all vertices $v$, the inequation to (at least) one of the outgoing edges of $v$ is sharp, and the edges for which they are sharp describe co-optimal strategies. This valuation of the game is thus the only valuation that defines strategies for both players, which shows the second claim.
Moreover, if $\sigma$ describes co-optimal strategies, then $f_{\sigma}(\mathsf{val})=0$ holds for $\mathsf{val}=\mathsf{val}_{\mathcal{G}}$ (and for this valuation only), which establishes the first claim.
The theorem above ensures that, in case the condition at Line 6 holds, the algorithm terminates and provides the value of the game that then allows us to infer optimal strategies of both players. Otherwise we have to improve the objective function and make another iteration of the while loop. At Line 7, $\mathsf{ChooseBetterStrategies}$ can be any procedure that, for $f_{\sigma}(\mathsf{val})\neq 0$, provides a pair of strategies $\sigma^{\prime}$ better than $\sigma$ as defined as follows.
{defi}
Better Strategy
A strategy $\sigma^{\prime}$ for both players is better than a strategy $\sigma$ if, and only if,
$\min_{\mathsf{val}^{\prime}}f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<\min_{\mathsf{val}}f_{\sigma}(\mathsf{val})$
Given a game $\mathcal{G}$, a strategy $\sigma^{\prime}$ and the corresponding objective function $f_{\sigma^{\prime}}$ computed by $\mathsf{ObjectiveFunction}(\mathcal{G},\sigma)$, the call $\mathsf{LinearProgramming}(H,f_{\sigma^{\prime}})$ computes the minimal value of $f_{\sigma^{\prime}}$. If this value is strictly lower than the minimal value for $f_{\sigma}$, computed by $\mathsf{LinearProgramming}(H,f_{\sigma})$, then $\sigma^{\prime}$ is better than $\sigma$.
While we discuss how to implement this key function in the next section, we observe here that the algorithm terminates with a correct result with any implementation that chooses a better objective function in each round: correctness is due to it only terminating when $\mathsf{val}$ defines strategies for both players, which implies (cf. Theorem 1) that $\mathsf{val}$ is the valuation of $\mathcal{G}$ ($\mathsf{val}=\mathsf{val}_{\mathcal{G}}$) and all strategies defined by $\mathsf{val}$ are co-optimal. Termination is obtained by a finite number of positional strategies: by Observation 4, the value of the objective function of all of them is non-negative, while the objective function of an optimal solution to co-optimal strategies is 00 (cf. Theorem 1), which meets the termination condition of Line 6 (cf. Observation 4).
Corollary 2.
Algorithm 3 always terminates with the correct value.
5. Choosing Better Strategies
In this section, we discuss sufficient criteria for finding a procedure that efficiently implements $\mathsf{ChooseBetterStrategies}$. For this, we make four observations described in the next subsections:
  1. (1)
    All local improvements can be applied. A strategy $\sigma^{\prime}$ is a local improvement to a strategy $\sigma$ if $f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$ holds for the current valuation $\mathsf{val}$ (Section 5.1).
  2. (2)
    If the current valuation $\mathsf{val}$ does not define a pair of strategies $\sigma$ for both players and has no local improvements, then a better strategy $\sigma^{\prime}$ can be found by applying only switches from and to edges that already have offset 00 (Section 5.2).
  3. (3)
    The improvement mentioned in the previous point can be found for special games – the sharp and improving games defined in Section 5.3 – by trying a single edge switch.
  4. (4)
    Games can almost surely be made sharp and improving by adding random noise that retains optimal strategies (Section 5.4).
Together, these four points provide efficient means for finding increasingly better strategies, and thus to find the co-optimal strategies and the valuation of the discounted payoff game.
As a small side observation, when using a simplex based technique to implement $\mathsf{LinearProgramming}$ at Line 5 of Algorithm 3, then the pivoting of the objective function from point (1) and the pivoting of the basis can be mixed (this will be discussed in Section 5.5).
5.1. Local Improvements
The simplest – and most common – case of creating better strategies $\sigma^{\prime}$ from a valuation for the objective $f_{\sigma}$ for a strategy $\sigma$ is to consider local improvements. Much like local improvements in strategy iteration approaches, local improvements consider, for each vertex $v$, a successor $v^{\prime}\neq\sigma(v)$, such that $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))<\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for the current valuation $\mathsf{val}$, which is optimal for the objective function $f_{\sigma}$.
To be more general, our approach does not necessarily require to select only local improvements, but it can work with global improvements, though we cannot see any practical use of choosing differently. For instance, if we treat the function as a global improvement approach, we can update the value for a vertex $v$ such that it increases by 1 and update the value of another vertex $v^{\prime}$ such that it decreases by 2. The overall value of the function will decrease, even if locally some components increased their value. Interestingly, this cannot be done with a strategy improvement approach, as it requires to always locally improve the value of each vertex when updating.
Lemma 3.
If $\mathsf{val}$ is an optimal valuation for the linear programming problem at Line 5 of Algorithm 3 and $f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, then $\sigma^{\prime}$ is better than $\sigma$.
Proof 5.1.
The valuation $\mathsf{val}$ is, being an optimal solution for the objective $f_{\sigma}$, a solution to the system of inequations $H$. For a solution $\mathsf{val}^{\prime}$ which is optimal for $f_{\sigma^{\prime}}$, we thus have $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})\leq f_{\sigma^{\prime}}(\mathsf{val})<f_{\sigma}(\mathsf{val})$, which implies that $\sigma^{\prime}$ is better than $\sigma$ accordingly to Definition 4.1 of better strategy.
5.2. No Local Improvements
The absence of local improvements means that, for all vertices $v\in V$ and all outgoing edges $(v,v^{\prime})\in E$, $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))\geq\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$.
We define for a valuation $\mathsf{val}$ optimal for $f_{\sigma}$ (like the $\mathsf{val}$ produced in Line 5 of Algorithm 3):
  • $S_{\mathsf{val}}^{\sigma}=\{(v,v^{\prime})\in E\mid\mathsf{offset}(\mathsf{val},(v,v^{\prime}))=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))\}$ as the set of stale edges; naturally, every vertex has at least one outgoing stale edge: the one defined by $\sigma$;
  • $E_{\mathsf{val}}=\{(v,v^{\prime})\in E\mid\mathsf{offset}(\mathsf{val},(v,v^{\prime}))=0\}$ as the set of edges, for which the inequation for $\mathsf{val}$ is sharp; in particular, all edges in the basis of $H$ that defines $\mathsf{val}$ are sharp (and stale); and
  • $E_{\mathsf{val}}^{\sigma}$ as any set of edges between $E_{\mathsf{val}}$ and $S_{\mathsf{val}}^{\sigma}$ (i.e. $E_{\mathsf{val}}\subseteq E_{\mathsf{val}}^{\sigma}\subseteq S_{\mathsf{val}}^{\sigma}$) such that $E_{\mathsf{val}}^{\sigma}$ contains an outgoing edge for every vertex $v\in V$; we are interested to deal with sets that retain the game property that every vertex has a successor, we can do that by adding (non-sharp) stale edges to $E_{\mathsf{val}}$.
Note that $S_{\mathsf{val}}^{\sigma}$ is such a set, hence, an adequate set is easy to identify. We might, however, be interested in keeping the set small, and choosing the edges defined by $E_{\mathsf{val}}$ plus one outgoing edge for every vertex $v$ that does not have an outgoing edge in $E_{\mathsf{val}}$. The most natural solution is to choose the edge $(v,\sigma(v))\in E_{\mathsf{val}}^{\sigma}$ defined by $\sigma$ for each such vertex $v$.
{obs}
If $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is a DPG and $\sigma$ a strategy for both players such that $\mathsf{val}$ is an optimal solution for the objective $f_{\sigma}$ to the system of inequations $H$, then $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E_{\mathsf{val}}^{\sigma},w,\lambda)$ is also a DPG.
This simply holds because every vertex $v\in V$ retains at least one outgoing transition.
Lemma 4.
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be a DPG, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. If $\mathsf{val}$ does not define strategies of both players, then there is a better strategy $\sigma^{\prime}$ such that, for all $v\in V$, $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$.
Proof 5.2.
By Observation 5.2, $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E_{\mathsf{val}}^{\sigma},w,\lambda)$ is a DPG. Let $\mathsf{val}^{\prime}$ be the valuation of $\mathcal{G}^{\prime}$, and $\sigma^{\prime}$ be the strategies for the two players defined by it.
If $\mathsf{val}^{\prime}$ is also a valuation of $\mathcal{G}$, then we are done. However, this need not be the case, as the system of inequations $H^{\prime}$ for $\mathcal{G}^{\prime}$ is smaller than the set of inequations $H$ for $\mathcal{G}$, so $\mathsf{val}^{\prime}$ might violate some of the inequations that are in $H$, but not in $H^{\prime}$. Given that $\mathsf{val}^{\prime}$ is a valuation for $\mathcal{G}^{\prime}$, it satisfies all inequations in $H^{\prime}$. Moreover, since $\mathsf{val}$ also satisfies all inequations of $H^{\prime}$, it follows that the same inequations hold for every convex combination of $\mathsf{val}$ and $\mathsf{val}^{\prime}$.
We now note that the inequations of $H$ that are not in $H^{\prime}$ are not sharp for $\mathsf{val}$. Thus, there is an $\varepsilon\in(0,1]$ such that the convex combination $\mathsf{val}_{\varepsilon}=\varepsilon\cdot\mathsf{val}^{\prime}+(1-\varepsilon)\cdot\mathsf{val}$ is a solution to those inequations.
We now have $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})>0$ and $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})=0$. For an optimal solution $\mathsf{val}^{\prime\prime}$ of $H$ for the objective $f_{\sigma^{\prime}}$, this provides $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})\leq f_{\sigma^{\prime}}(\mathsf{val}_{\varepsilon})<f_{\sigma}(\mathsf{val})$.
Therefore $\sigma^{\prime}$ is better than $\sigma$.
When using the most natural choice, $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$, this translates in keeping all transitions, for which the offset is not 00, while changing some of those, for which the offset already is 00. This is a slightly surprising choice, since to progress one has to improve on the transitions whose offset is positive, and ignore those with offset 0.
5.3. Games with Efficient Objective Improvement
In this subsection, we consider sufficient conditions for finding better strategies efficiently. Note that we only have to consider cases where the termination condition (Line 6 of Algorithm 3) is not met.
The simplest condition for efficiently finding better strategies is the existence of local improvements. In particular, it is easy to find, for a given valuation $\mathsf{val}$, strategies $\sigma^{\prime}$ for both players such that $f_{\sigma^{\prime}}(\mathsf{val})\leq f_{\sigma^{\prime\prime}}(\mathsf{val})$ holds for all strategies $\sigma^{\prime\prime}$. When there are local improvements, we can obtain a better strategy simply by applying them. This leaves the case in which there are no local improvements, but where $\mathsf{val}$ also does not define strategies for the two players. We have seen that we can obtain a better strategy by only swapping edges, for which the inequations are sharp (Lemma 4).
We now describe two conditions that, when simultaneously met, allow us to efficiently find better strategies: that games are sharp and improving.
Sharp games.
To do this efficiently, it helps if there are always $|V|$ inequations that are sharp for a valuation $\mathsf{val}$ that minimises $f_{\sigma}(\mathsf{val})$ for some $\sigma$. Of course, if $\mathsf{val}$ is a solution returned by the simplex method, then there must be at least $|V|$ sharp inequations, as $\mathsf{val}$ is the solution defined by making $|V|$ inequations sharp (namely those that form the basis). By requiring that there are exactly $|V|$ sharp inequations we require that an optimal valuation defines a basis. We call such a set of inequations $H$ and games that define them sharp games.
{defi}
Sharp Game
A game $\mathcal{G}$ is called sharp for the valuation $\mathsf{val}$ if, and only if, $\mathsf{val}$ satisfies exactly $|V|$ inequations $H$ computed by $\mathsf{Inequations}(\mathcal{G})$.
Improving games.
The second condition, which allows us to identify better strategies efficiently, is to assume that, for every strategy $\sigma$ for both players, if a valuation $\mathsf{val}$ defined by a basis is not optimal for $f_{\sigma}$ under the constraints $H$, then there is a single basis change that improves it. We call such games improving. Being improving is the property we really need, while sharpness is a property that allows us to almost surely make a game improving by applying small noise on the objective functions.
DPGs are not always improving. Being improving is a very useful property as it removes the problem that the simplex method can stall, as a single basis change is enough for improving games to improve the value of the objective function (hence the name). For non-improving games, adjacent solutions might not improve although the current basis does not define a solution that is optimal for the current objective function.
{defi}
Improving Game
A game $\mathcal{G}$ is called $\sigma$-improving for an objective function $f_{\sigma}$ defined by a joined strategy $\sigma$ and a non-optimal valuation $\mathsf{val}$ defined by a basis for $f_{\sigma}$ if, and only if, there is a single basis change that leads to a better strategy.
A game $\mathcal{G}$ is called improving if it is improving for all objective functions $f_{\sigma}$ defined by a joined strategy $\sigma$ and all valuations $\mathsf{val}$ defined by bases that are not optimal w.r.t. $f_{\sigma}$.
We call a valuation $\mathsf{val}^{\prime}$ whose basis can be obtained from that of $\mathsf{val}$ by a single change to the basis of $\mathsf{val}$ a neighbouring valuation to $\mathsf{val}$. We show that, for improving games, we can refine the result of Lemma 4 so that the better strategy $\sigma^{\prime}$ also guarantees $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ for some neighbouring valuation $\mathsf{val}^{\prime}$ to $\mathsf{val}$.
This allows us to consider $O(|E|)$ basis changes and, where they define a valuation, to seek optimal strategies for a given valuation. Finding an optimal strategy for a given valuation is straightforward.
Theorem 5.
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for $f_{\sigma}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is (at least) one neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
Such a strategy $\sigma^{\prime}$ is better than $\sigma$, and it can be selected in a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
Proof 5.3.
We apply Lemma 4 for $E_{\mathsf{val}}^{\sigma}$ that, we recall, corresponds to $E_{\mathsf{val}}\cup\{(v,\sigma(v)\mid v\in V\}$, and use the resulting better strategy $\sigma^{\prime}$ from this set $E^{\sigma}_{\mathsf{val}}$. Note that this means $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$, as there are no local improvements.
We now consider an arbitrary sequence of evaluations $\mathsf{val}_{0},\mathsf{val}_{1},\ldots,\mathsf{val}_{n}$ obtained by calls to the simplex algorithm on the linear program defined by the constraints $H$ obtained from $\mathcal{G}$. The evaluations correspond to adjacent corners of the polytope on which the simplex algorithm moves. The sequence starts from $\mathsf{val}_{0}=\mathsf{val}$ and ends with $\mathsf{val}_{n}=\mathsf{val}^{\prime}$. The value of the new objective function $f_{\sigma^{\prime}}$ can only decrease along this path between the corners of the polytope and $\mathsf{val}^{\prime}$ minimises $f_{\sigma^{\prime}}$.
Note that, since $\sigma^{\prime}$ is better than $\sigma$ by Lemma 4, we have that $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$. Moreover, the sharpness of $\mathcal{G}$ implies that $\mathsf{val}_{1}\neq\mathsf{val}_{0}$ and, since $\mathcal{G}$ is improving (and therefore $\sigma^{\prime}$-improving), this provides that $f_{\sigma^{\prime}}(\mathsf{val}_{1})<f_{\sigma^{\prime}}(\mathsf{val}_{0})$. Finally, we have already established $f_{\sigma^{\prime}}(\mathsf{val}_{0})=f_{\sigma}(\mathsf{val}_{0})$.
Thus, when only applying a single basis change, we move to a fresh value $\mathsf{val}^{\prime\prime}=\mathsf{val}_{1}$ such that $f_{\sigma^{\prime}}(\mathsf{val}_{1})<f_{\sigma}(\mathsf{val})$ for some $\sigma^{\prime}$.
Note that $\sigma^{\prime}$ was supplied by Lemma 4, so that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds.
While we can restrict the selection of $\sigma^{\prime}$ to the strategies that comply with the restriction $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, there is no particular reason for doing so; as soon as we have a neighbouring valuation $\mathsf{val}^{\prime}$, we can identify a pair of strategies $\sigma^{\prime}$ for which $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})$ is minimal, and select $\sigma^{\prime}$ if $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$ holds.
5.4. Making Games Sharp and Improving
Naturally, not every game is improving, or even sharp. In this subsection, we first discuss how to almost surely make games sharp by adding sufficiently small random noise to the edge weights, and then discuss how to treat games that are not improving by assigning randomly chosen factors, with which the offsets of edges are weighted. Note that these are ‘global’ adjustments of the game that only need to be applied once, as it is the game that becomes sharp and improving, respectively.
Starting with the small noise to add on the edge weights, we first create notation for expressing how much we can change edge weights, such that joint co-optimal strategies of the resulting game are joint co-optimal strategies in the original game. To this end, we define the gap of a game.
Gap of a game.
Before defining the gap of the game we recall that $\mathsf{val}_{\sigma}$ is the valuation of the joint strategy $\sigma$ (while $\mathsf{val}$ denotes the outcome of the optimisation problem). We use this valuation to argue that a small distortion of the edge weights cannot turn non-co-optimal strategies into co-optimal ones, and we define the gap of the game to reason about how small is small enough to retain non-co-optimality. The value of a joint strategy itself is useful for this, because it is much easier to access than the result of optimisation. As a valuation of a non-co-optimal joint strategy is no solution of a game, we will encounter negative offsets: they occur for exactly those edges, for which the associated inequation is not satisfied by $\mathsf{val}_{\sigma}$.
{defi}
Contraction and Gap of a Game
Given a game $\mathcal{G}$, we call its contraction $\lambda^{*}=\max\{\lambda_{e}\mid e\in E\}$. For a non-co-optimal joint strategy $\sigma$, we call the gap of $\sigma$ for $\mathcal{G}$ the value $\gamma_{\sigma}=-\min\{\mathsf{offset}(\mathsf{val}_{\sigma},e)\mid e\in E\}$. We call the gap of $\mathcal{G}$ the value $\gamma=\min\{\gamma_{\sigma}\mid\sigma$ is a non-co-optimal joint strategy}}\}}, i.e. the minimal such value.
Note that $\gamma_{\sigma}>0$ always holds for non-co-optimal strategies; for the minimal [^3] This skips over the case where all strategies are co-optimal, but that case is trivial to check and such games are trivial to solve, so that we ignore this case in this subsection. value $\gamma$, $\gamma>0$ thus always holds. Therefore, it is possible to obtain a better value by playing a different edge than the one selected by $\sigma$.
We now use the gap of a game $\gamma$ to define the magnitude of a change to all weights, such that all strategies that used to have a gap still have one, and thus that all co-optimal joint strategies from the distorted game are also co-optimal joint strategies in the undistorted game.
Lemma 6.
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be a DPG with contraction $\lambda^{*}$ and gap $\gamma$, and let
$\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ differ from $\mathcal{G}$ only in the edge weights such that, for all $e\in E$, $|w_{e}-w_{e}^{\prime}|<\frac{1-\lambda^{*}}{2}\gamma$ holds. Then a joint co-optimal strategy from $\mathcal{G}^{\prime}$ is also co-optimal for $\mathcal{G}$.
Proof 5.4.
We argue that the small weight disturbance, $|w_{e}-w_{e}^{\prime}|<\frac{1-\lambda^{*}}{2}\gamma$ for all $e\in E$, provides a small difference in the valuation that does not affect co-optimality. By using the definition of $\mathsf{val}_{\sigma}$ and the triangle inequality we can estimate the difference between the valuations. Precisely, for all joint strategies $\sigma$, we have for $\mathsf{val}$ on $\mathcal{G}$, and $\mathsf{val}^{\prime}$ on $\mathcal{G}^{\prime}$, that $|\mathsf{val}_{\sigma}(v)-\mathsf{val}_{\sigma}^{\prime}(v)|<\frac{1}{1-\lambda^{*}}\frac{1-\lambda^{*}}{2}\gamma=\frac{\gamma}{2}$. In fact, $\sigma$ defines a play $\rho=e_{0}e_{1}e_{2}\ldots$, and we have $\mathsf{val}_{\sigma}(v)=\mathsf{out}(\rho)=\sum_{i=0}^{\infty}w_{e_{i}}\prod_{j=0}^{i-1}\lambda_{e_{j}}$ and $\mathsf{val}_{\sigma}^{\prime}(v)=\sum_{i=0}^{\infty}w_{e_{i}}^{\prime}\prod_{j=0}^{i-1}\lambda_{e_{j}}$. This provides
$|\mathsf{val}_{\sigma}(v)-\mathsf{val}_{\sigma}^{\prime}(v)|\leq\sum_{i=0}^{\infty}|w_{e_{i}}-w_{e_{i}}^{\prime}|\prod_{j=0}^{i-1}\lambda_{e_{j}}<\frac{1-\lambda^{*}}{2}\gamma\sum_{i=0}^{\infty}\prod_{j=0}^{i-1}\lambda_{e_{j}}\leq\frac{1-\lambda^{*}}{2}\gamma\sum_{i=0}^{\infty}(\lambda^{*})^{i}=\frac{\gamma}{2}.$
If we assume that $\sigma$ is not co-optimal for $\mathcal{G}$, then we have an edge $e=(v,v^{\prime})$ with $-\mathsf{offset}(\mathsf{val}_{\sigma},e)=\gamma_{\sigma}\geq\gamma$. Using $\mathsf{offset}^{\prime}$ to indicate the use of $w_{e}^{\prime}$ for $\mathcal{G}^{\prime}$, by applying the triangle inequality we get
$\begin{array}[]{cl}|\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)-\mathsf{offset}(\mathsf{val}_{\sigma},e)|&\leq|\mathsf{val}_{\sigma}^{\prime}(v)-\mathsf{val}_{\sigma}(v)|+|w_{e}-w_{e}^{\prime}|+\lambda_{e}|\mathsf{val}_{\sigma}(v^{\prime})-\mathsf{val}_{\sigma}^{\prime}(v^{\prime})|\\ &<\frac{\gamma}{2}+\frac{1-\lambda^{*}}{2}\gamma+\lambda_{e}\frac{\gamma}{2}\leq\gamma\\ \end{array}$
Together with the fact that $-\mathsf{offset}(\mathsf{val}_{\sigma},e)\geq\gamma$, this provides $\mathsf{offset}^{\prime}(\mathsf{val}_{\sigma}^{\prime},e)<0$.
Thus, $\sigma$ is not co-optimal for $\mathcal{G}^{\prime}$ either.
Lemma 7.
Given a DPGs $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$, the DPG $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ resulting from $\mathcal{G}$ by adding independently uniformly at random drawn values from an interval $(-\varepsilon,\varepsilon)$ to every edge weight, will almost surely result in a sharp game.
Note that these distributions are continuous. They could be replaced by any other distribution over these intervals that has 00 weight for all individual points.
Proof 5.5.
In this proof, we show a stronger property: every two different bases almost surely define different valuations (regardless of whether or not these valuations are solutions). There are only finitely many bases, and it suffices to compare two arbitrary but fixed ones, say $b_{1}$ and $b_{2}$.
What we show is that it happens with probability 00 that $b_{1}$ and $b_{2}$ are bases and define the same solution. If $b_{1}$ or $b_{2}$ is not a basis, we are done, too.
As $b_{1}$ and $b_{2}$ are different, there will be one edge $e=(v,v^{\prime})$ that occurs in $b_{1}$, but not in $b_{2}$. As all weight disturbances are drawn independently, we assume without loss of generality that the weight disturbance to this edge is drawn last.
Now, the valuation $\mathsf{val}_{1}$ defined by $b_{1}$ does not depend on this final draw. For $\mathsf{val}_{1}$, there is a value $w_{e}^{\prime}=\mathsf{val}_{1}(v)-\lambda_{e}\mathsf{val}_{1}(v^{\prime})$ that defines the weight $w_{e}^{\prime}$ that $e$ would need to have, in order to make the inequation sharp for $e$.
For the valuation $\mathsf{val}_{1}$ defined by $b_{1}$ to be equal to the valuation $\mathsf{val}_{2}$ defined by $b_{2}$, the weight for the edge $e$ (after adding the drawn distortion) needs to be exactly $w_{e}^{\prime}$. (This is a necessary condition, but not necessarily a sufficient condition.) There is at most one value for the disturbance that would provide for this, and this disturbance for the weight of $e$ is sampled with a likelihood of 00.
Putting these two results together, we get:
Corollary 8.
Given a pair of DPGs $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ with contraction $\lambda^{*}$ and gap $\gamma$, and $\mathcal{G}^{\prime}=(V_{\min},V_{\max},E,w^{\prime},\lambda)$ obtained from $\mathcal{G}$ by adding independently uniformly at random drawn values from an interval $(-\varepsilon,\varepsilon)$ to every edge weight, for some $\varepsilon\leq\frac{1-\lambda^{*}}{2}\gamma$, then, a joint co-optimal strategy from $\mathcal{G}^{\prime}$ is also co-optimal for $\mathcal{G}$, and $\mathcal{G}^{\prime}$ is almost surely sharp.
Note that we can estimate the gap cheaply when all coefficients in $\mathcal{G}$ are rational. The gap is defined as the minimum of $\gamma_{\sigma}$ over non-co-optimal joint strategies $\sigma$, and we start by fixing such a joint strategy.
For a given $v\in V$, $\sigma$ defines a run $\rho=e_{0}e_{1}e_{2}\ldots$ in the form of a “lasso path”, which consists of a (possibly empty) initial path $e_{0},\ldots,e_{k}$, followed by an infinitely often repeated cycle $e_{0}^{\prime},\ldots,e_{\ell}^{\prime}$, where the only vertex that occurs twice on the path $e_{0},\ldots,e_{k},e_{0}^{\prime},\ldots,e_{\ell}^{\prime}$ is the vertex reached before $e_{0}^{\prime}$ and after $e_{\ell}^{\prime}$, while all other vertices occur only once.
In this run, an edge $e_{i}$ occurs only once and contributes $\prod_{j=0}^{i-1}\lambda_{e_{j}}w_{e_{i}}$ to the value of this run, while an edge $e_{i}^{\prime}$ occurs infinitely often and contributes the value $\frac{(\prod_{j=0}^{k}\lambda_{e_{j}})\cdot(\prod_{j=0}^{i-1}\lambda_{e_{j}^{\prime}})\cdot w_{e_{i}^{\prime}}}{1-\prod_{j=0}^{\ell}\lambda_{e_{j}^{\prime}}}$ to the value of the run. Now all we need to do is to estimate a common denominator.
To do this, let $\mathsf{denom}(r)$ be the denominator of a rational number $r$. It is easy to see that
$\mathsf{common^{\prime}}=\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})\cdot\mathsf{denom}(w_{(v,\sigma(v))})$
is a common denominator of all expressions of the first type ($\prod_{j=0}^{i-1}\lambda_{e_{j}}w_{e_{i}}$ and of the nominator from the second ($(\prod_{j=0}^{k}\lambda_{e_{j}})\cdot(\prod_{j=0}^{i-1}\lambda_{e_{j}^{\prime}})\cdot w_{e_{i}^{\prime}}$).
This leaves the contribution of the denominator of the fraction from the second form, which is the nominator of the terms $1-\prod_{j=0}^{\ell}\lambda_{e_{j}^{\prime}}$. But as the term is between 00 and $1$, its value is strictly smaller than the value of its denominator; we thus have
$\mathsf{nom}(1-\prod_{j=0}^{\ell}\lambda_{e_{j}^{\prime}})\leq\prod_{j}^{l}\mathsf{denom}(\lambda_{e_{j}^{\prime}}).$
Considering that the cycles do not share states (or edges), we in particular have that smallest common multiple $\mathsf{scp}$ of these nominators is smaller than the product of all relevant edge denominators; we in particular have
$\mathsf{scp}\leq\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})\ ,$
such that
$\mathsf{common}=\mathsf{common^{\prime}}\cdot\mathsf{scp}<\prod\limits_{v\in V}\mathsf{denom}(\lambda_{(v,\sigma(v))})^{2}\cdot\mathsf{denom}(w_{(v,\sigma(v))})$
is a common denominator for all occurring values.
Looking at an edge $e$ that defines $\gamma_{\sigma}$, then $\gamma_{\sigma}$ can be written with the denominator
$\mathsf{common}\cdot\mathsf{denom}(\lambda_{e}w_{e})\ .$
Obviously, the nominator is at least $1$.
Estimating $\gamma$ can thus be done by using the highest possible denominators available in $\mathcal{G}$. The representation of the resulting estimate $\gamma$ is polynomial in the size of $\mathcal{G}$.
Biased sum of offsets.
We modify sharp games that are not improving. This can be done by redefining the function $\mathsf{offset}$ as follows:
$\mathsf{offset}^{\prime}(\mathsf{val},(v,v^{\prime}))=\alpha_{(v,v^{\prime})}\cdot\mathsf{offset}(\mathsf{val},(v,v^{\prime}))$
Such $\mathsf{offset}^{\prime}$ is defined for every edge $e=(v,v^{\prime})$, where each $\alpha_{e}$ is a positive number, which we call the offset factor. Based on this change, we re-define all offset definitions and the objective function with a primed version that uses these $\mathsf{offset}^{\prime}$ factors, and thus also what an improving DPG is.
Theorem 9.
Let $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ be an improving DPG for a given set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$, $\sigma$ a strategy for both players, $\mathsf{val}$ an optimal solution returned at Line 5 of Algorithm 3 for the adjusted function $f_{\sigma}^{\prime}$, and let there be no local improvements of $\sigma$ for $\mathsf{val}$. Then there is neighbouring valuation $\mathsf{val}^{\prime\prime}$ to $\mathsf{val}$ such that there is a better strategy $\sigma^{\prime}$ that satisfies $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})<f_{\sigma}(\mathsf{val})$.
Such a strategy $\sigma^{\prime}$ is better than $\sigma$, and it can be selected in a way that $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$ holds for all $v\in V$ for a given set $E_{\mathsf{val}}^{\sigma}$.
The proof from Theorem 5 can be repeated verbatim with the new offset definition.
Theorem 10.
If each offset factor $\alpha_{e}$ is selected independently uniformly at random from a bounded interval of positive numbers444or from any other distribution over positive numbers that has 00 weight for all individual points, then a sharp DPG $\mathcal{G}=(V_{\min},V_{\max},E,w,\lambda)$ is almost surely improving for a sampled set of positive numbers $\{\alpha_{e}>0\mid e\in E\}$.
Proof 5.6.
We show the claim for two arbitrary but fixed valuations $\mathsf{val}_{1}$ and $\mathsf{val}_{2}$ defined by two different bases, $b_{1}$ and $b_{2}$, respectively, that satisfy the inequations in $H$, and an arbitrary but fixed joint strategy $\sigma$. As there are finitely many bases and finitely many joint strategies, satisfying the requirement almost surely for them entails that the requirement is satisfied almost surely for the game.
As $\mathcal{G}$ is sharp, we have $\mathsf{val}_{1}\neq\mathsf{val}_{2}$.
We first assume for contradiction that $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))=\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ holds for all $v\in V$. We pick a vertex $v$ such that $|\mathsf{val}_{1}(v)-\mathsf{val}_{2}(v)|>0$ is maximal; such a vertex exists since $\mathsf{val}_{1}\neq\mathsf{val}_{2}$. For this $v$, $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))=\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ entails $\mathsf{val}_{1}(v)-\mathsf{val}_{2}(v)=\lambda_{(v,\sigma(v))}(\mathsf{val}_{1}(\sigma(v))-\mathsf{val}_{2}(\sigma(v)))$. Using $\lambda_{e}\in[0,1)$ for $e=(v,\sigma(v))$, we get $|\mathsf{val}_{1}(\sigma(v))-\mathsf{val}_{2}(\sigma(v))|>|\mathsf{val}_{1}(v)-\mathsf{val}_{2}(v)|>0$, which contradicts the maximality of $v$.
We therefore have that $\mathsf{offset}(\mathsf{val}_{1},(v,\sigma(v)))\neq\mathsf{offset}(\mathsf{val}_{2},(v,\sigma(v)))$ holds for some $v\in V$. As the $\alpha_{e}$ are drawn independently, we can assume w.l.o.g. that $\alpha_{(v,\sigma(v))}$ is drawn last. There is at most one value $\alpha_{(v,\sigma(v))}^{\prime}$ for which the following condition is not satisfied:
$\sum_{v\in V}\mathsf{offset}^{\prime}(\mathsf{val}_{1},(v,\sigma(v)))\neq\sum_{v\in V}\mathsf{offset}^{\prime}(\mathsf{val}_{2},(v,\sigma(v))).$
It therefore almost surely holds for all strategies that all basis-induced valuations have a pairwise distinct image by the objective function associated to the strategy. This immediately implies that the game is improving.
First, it is sharp by assumption (the factors $\alpha_{e}$ do not affect sharpness, as they only affect the objective functions $f_{\sigma}$). Second, assume that a basis defines a solution $\mathsf{val}$ that is not optimal w.r.t. the adjusted $f_{\sigma}$, then the simplex algorithm will pick a neighbouring basis that defines a solution $\mathsf{val}^{\prime}$ with $f_{\sigma}(\mathsf{val}^{\prime})\leq f_{\sigma}(\mathsf{val})$. Finally, as we have just shown that $f_{\sigma}(\mathsf{val}^{\prime})\neq f_{\sigma}(\mathsf{val})$ holds almost surely and consider one of the games where this holds, this implies $f_{\sigma}(\mathsf{val}^{\prime})<f_{\sigma}(\mathsf{val})$, which is the definition of improving for sharp games.
Thus, we can almost surely obtain sharpness by adding small noise to the weights, and almost surely make games improving by considering the offsets of the individual edges with a randomly chosen positive weight. This guarantees cheap progress for the case where there are no local improvements.
5.5. Mixing Pivoting on the Simplex and of the Objective
When using a simplex based technique to implement $\mathsf{LinearProgramming}$ (Line 5 of Algorithm 3), then the algorithm mixes three approaches that stepwise reduce the value of $f_{\sigma}(\mathsf{val})$:
  1. (1)
    The simplex algorithm updates the basis, changing $\mathsf{val}$ (while retaining the objective function $f_{\sigma}$).
  2. (2)
    Local updates, that change the objective function $f_{\sigma}$ (through updating $\sigma$) and retain $\mathsf{val}$.
  3. (3)
    Non-local updates.
Non-local updates are more complex than the other two, and the correctness proofs make use of the non-existence of the other two improvements. For both reasons, it seems natural to take non-local updates as a last resort.
The other two updates, however, can be freely mixed, as they both bring down the value of $f_{\sigma}(\mathsf{val})$ by applying local changes. That the improvements from (1) are given preference in the algorithm is a choice made to keep the implementation of the algorithm for using linear programs open, allowing, for example, to use ellipsoid methods [Kha79] or inner point methods [Kar84] to keep this step tractable.
6. Experimental Evaluation
To evaluate the performance of the novel approach, we have implemented it in an experimental framework in C++. Together with the objective improvement algorithm (OI, for short) we have also implemented the classic asymmetric variant of the strategy improvement algorithm (SI, for short) that has been described in Section 3. Note that our implementation uses floating point numbers with double precision (type doubles), while the approach itself assumes infinite precision. As a consequence, match checks are not possible and the solutions can be affected by an $\epsilon$ error due to the precision of the library (that comes from the external linear programming solver too). For most parts, this could be rectified by using fractions. However, while fractions with infinite precision are available with the right libraries, their use would lead to an overhead and would not solve the precision problem of the external linear programming solver. We also note that the small noise added in Section 5.4 to almost surely make a game sharp and improving requires a continuous value space.
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 00000555510101010151515152020202025252525303030303535353540404040Game sizeNumber of iterations (avg)SIOI
Figure 2. Comparisons of the average number of iterations (LP calls) on games with two successors for each vertex.
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 00045454545100100100100200200200200300300300300400400400400500500500500600600600600700700700700Game sizeNumber of updates (avg)SIOI
Figure 3. Comparisons of the average number of local strategy updates for games with two successors per vertex.
Evaluation comes with a number of difficulties. First, there are no actual concrete benchmarks for DPGs. We have therefore translated parity games that model synthesis problems into DPGs. The translation process is exponential in the number of priorities of the parity games. For this reason we could translate and test only two types of concrete games: the Elevator and Language Inclusion. Besides these concrete games, we have also generated random games of various sizes. We have considered games with rational weights in the interval $[-250,250]$ on the edges, as games with integer weights represent a restricted class of games. The range of rational weights is not crucial as it suffice to multiply their values to scale the interval up or down. Moreover, we use a uniform discount factor of $0.9$, since for games in which each vertex has different weights on the edges, non-uniform discount factors do not make games any easier or harder to solve. Using a low discount factor (e.g. 0.5 or below) make, instead, games easier to solve, because with small discount factors the solution (i.e. the valuation $\mathsf{val}$) of the vertices converges after few steps, while with a discount factor close to 1 the valuation requires many steps to converge. Finally, it is worth noting that games with such random weights on the edges are also almost surely sharp, hence, it has not been necessarily to implement the sharpening technique described in Section 5.4.
The second obstacle is that, for an implementation, one has to choose how to update the current “state”, both for SI and OI. For SI, we use a greedy all switch rule: we update all local decisions where there is a local improvement available, and where there are multiple, we choose one with the maximal local improvement. For OI, there are more choices to make. As mentioned in Section 5.5, in simplex-based approaches, we can freely mix (1) the standard updates of the basis for a fixed objective function, and (2) updates of the objective function itself, and in most steps, both options are available. For (1), there are a plethora of update rules for the simplex, and for (2), there is also a range of ways of how to select the update. We have dodged this question by not directly implementing a solver, but by, instead, solving the linear programs with an off-the-shelf solver. We have used the LP solver provided by the ALGLIB library [^5] Library libalglib version 3.16.0 for both OI and SI. This does not restrict us to the use of simplex-based solvers, while it would restrict simplex-based solutions to always giving priority to (1) over (2). Where the solver does provide a solution for a given objective function, we use rule (2) with a greedy all switch rule.
One cost of doing this is that we do not have a similarly straightforward integration of (3) non-local updates, where we are in the optimal solution for an objective function that cannot be improved with rule (2). Instead of using the method from Theorem 5 to slightly adjust a simplex method to find an improvement for improving games (knowing that the games we create are almost surely improving), we use the general rule from Lemma 4. In principle, this might lead to large plateaux that are hard to exit, but we considered this very unlikely, at least for the improving games we create, and our experiments have confirmed this.
This setup allows us to fairly compare OI with SI by counting the number of linear program instances that need to be solved on the way.
As our first benchmark set we selected games with two outgoing edges per vertex. This type of games have a relatively small strategy space, which should make them a relatively easy target for SI in that SI should require fewer steps to identify the optimal strategy than OI. Figures 2 and 3 show the number of iterations and strategy updates required to solve a game of size ranging from 100 to 1000 vertices. We selected 1000 games, each of the ten points in the graph shows the mean value for a cluster of 100 games of the corresponding size.
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 000001111222233334444555566667777888899991010101011111111Game sizeNumber of iterations (avg)SIOI
Figure 4. Comparisons of the average number of iterations (LP calls) for games with many (5 to 10) successors per vertex.
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 0001601601601603003003003006006006006009009009009001,20012001{,}2001 , 2001,50015001{,}5001 , 5001,80018001{,}8001 , 8002,10021002{,}1002 , 100Game sizeNumber of updates (avg)SIOI
Figure 5. Comparisons of the average number of local strategy updates for games with many (5 to 10) successors per vertex.
The number of iterations (Figure 2) represents how many times the linear programming solver has been called. As expected, SI is more efficient than OI at solving these games in terms of LP calls. What was less expected is how the advantage of SI over OI grows with the size of the game. Solving the linear programs, instead, is simpler for OI, at least when using a simplex style algorithm [^6] Simplex style algorithms (reported in Algorithm 1) operate in two phases: in the first phase, they look for any basis that defines a solution to the constraint system (Line 1), and in the second, they optimise (while loop). For OI, the first phase can be skipped for all but the first call; this is because the constraint system does not change, so that the previous solution can be used as a starting point. For SI, the constraint system changes; here, the previous solution is never a solution to the new linear program.. We also note that only 55 (5.5%) of the 1000 games considered overall reached a non-improving state. In these rare cases, this has been solved by choosing a stale switch (i.e. a switch of strategy that leads to the same valuation value).
The number of updates, instead, counts how many times each vertex changed its strategy before the next call to the LP solver, and then provides the sum of these switches. If the ownership of the vertices of a game is not balanced (e.g. one player owns 80% of the vertices), most of the solving effort is charged to the LP solver, while the SI algorithm itself needs to update very few strategies. On the contrary, OI updates the strategies for both the players. It would therefore stand to reason to expect that the number of local strategy updates would give an extra advantage to SI. This is, however, not the case: Figure 3 shows that, while OI needs to update the strategy for at least twice as many vertices, it needs only around a third more local updates for this.
In the second benchmark set we selected games with more moves per vertex, in the range $[5,10]$. In this case, the space of strategies is much wider than one of the games of the first benchmark set, and algorithms based on SI find these games harder to solve. Figures 4 and 5 show the comparison of OI and SI on the second benchmark set of 1000 games.
In Figure 4 we can observe a slowly but consistently growing gap in the number of LP calls (iterations). In addition to the advantage of simpler LP problems, we see that in these slightly more complex games OI also requires fewer calls, with SI needing some 2.5 to 3 times as many. For this benchmark set, only 7 (0.7%) of the games reached a non-improving state at any point.
Interestingly, Figure 5, instead, depicts a symmetric behaviour of the graph of Figure 3, but this time with the lanes of the two solvers swapped. Thus, although OI considers the strategy at all vertices (instead of at most half), the sum of local improvements is also slightly smaller.
When considering games with many transitions, namely games where the outdegree of every vertex is 10% of the number of vertices, the advantage of OI over SI grows further. Figure 6 shows a larger gap in the number of LP calls than Figure 4, and Figure 7 shows not only an advantage of OI over SI, but it also appears that the number of local strategy updates grows linearly for OI, but faster for SI. Only 3 (0.3%) of the games reached a non-improving state at some point of the solving process.
Finally, for the concrete problems translated from parity games, all these games turned out to be easy to solve so that a single LP call suffices to find the optimal valuation. Therefore, in Table 2, instead of showing the number of iterations and updates, we show the solution time in seconds that provides a measure of the complexity of the game in terms of size of the linear system. Most of the games can be solved in less than one second. For the Elevator family of games, their size grows rapidly, so that we could not translate more than the first 5 instances. For the Language Inclusion games, instead, we report the games solved within 15 minutes.
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , 000001111222233334444555566667777888899991010101011111111121212121313131314141414Game sizeNumber of iterations (avg)SIOI
Figure 6. Comparisons of the average number of iterations (LP calls) for games with a linear number of successors per vertex (10%).
1001001001002002002002003003003003004004004004005005005005006006006006007007007007008008008008009009009009001,00010001{,}0001 , $0001{,}$ 0001,50015001{,}5001 , 5002,00020002{,}0002 , 0002,50025002{,}5002 , 5003,00030003{,}0003 , 0003,50035003{,}5003 , 5004,00040004{,}0004 , 0004,50045004{,}5004 , 5005,00050005{,}0005 , 0005,50055005{,}5005 , 5006,00060006{,}0006 , 0006,50065006{,}5006 , 5007,00070007{,}0007 ,  [^0] Game sizeNumber of updates (avg)SIOI
Figure 7. Comparisons of the average number of local strategy updates for games with a linear number of successors per vertex (10%).
Benchmark Positions Moves Time
Elevator 1 $36$ $54$ 00
Elevator 2 $144$ $234$ 00
Elevator 3 $564$ $950$ 00
Elevator 4 $2688$ $4544$ $2$
Elevator 5 $15683$ $26354$ $16$
Language Inclusion 1 $170$ $1094$ 00
Language Inclusion 2 $304$ $1222$ 00
Language Inclusion 3 $428$ $878$ 00
Language Inclusion 4 $628$ $1538$ 00
Language Inclusion 5 $509$ $2126$ 00
Language Inclusion 6 $835$ $2914$ 00
Language Inclusion 7 $1658$ $4544$ $1$
Language Inclusion 8 $14578$ $17278$ $47$
Language Inclusion 9 $25838$ $29438$ $279$
Table 2. Experiments on concrete verification problems.
7. Discussion
There is widespread belief that mean payoff and discounted payoff games have two types of algorithmic solutions: value iteration [FGO20, Koz21] and strategy improvement [Lud95, Pur95, BV07, Sch08, STV15]. We have added a third method, which is structurally different and opens a new class of algorithms to attack these games. Moreover, our new symmetric approach has the same proximity to linear programming as strategy improvement algorithms, which is an indicator of efficiency.
The lack of a benchmarking framework for existing algorithms prevents us from testing and comparing an eventual implementation thoroughly, but we have compared OI against SI on random games and synthesis parity translated games in Section 6. To keep this comparison fair (and simple), we have put as much as possible of both problems into LP calls, using greedy all switch updates in both cases. We found that SI performs better on random games with very few (two) successors per vertex, while OI already comes out on top with few (five to ten), and shines with many (10% of the vertices) successors. This was a bit unexpected to us, as we thought that such a naive implementation of the new concept could not possibly compete.
Our results for random games with a low (five to ten) outdegree suggests that the number of local updates seen grows linearly with the size of the game for both SI and OI, and that the number of iterations initially grows and later plateaus. This may well be the same for games with a minimal outdegree of two, and the data clearly supports this for SI, but the number of LP calls growth almost linearly for OI. This might be an outlier in the behaviour and it is well possible that OI plateaus later, as it is unlikely that games with a low and very low outdegree behave fundamentally different. For SI, this was entirely expected, and it is maybe not overly surprising for OI either. Generally speaking, OI appears to do better than SI measured in the number of LP calls on random games, except where the outdegree is tiny.
Naturally, a fresh approach opens the door to much follow-up research. A first target for such research is the questions on how to arrange the selection of better strategies to obtain fewer updates, either proven on benchmarks, or theoretically in worst, average, or smoothed analysis. Our implementation effectively updates the objective function only once an optimal solution for the current objective function is found. This is for two pragmatic reasons: it allows us to use a solver for linear programs as a black box, and we feel that it provides the best comparison with strategy improvement. In terms of the three update rules stated in Section 5.5, this approach used (1) basis change where available, then (2) update of the objective function through updating the joint strategy, before turning to (3) non-local updates as a last resort.
Seeing that (1) and (2) are equally cheap and simple, we could also favour (2) over (1): where we have a solution $\mathsf{val}$ defined by a basis, we can first find a joint strategy $\sigma$ such that $f_{\sigma}(\mathsf{val})$ is minimal among all strategies, before turning to (1) and (3). This would also make sure that no vertex is ever visited twice. This could happen in principle when favouring (1) over (2), as a step that is good for one objective function might not be good for its successor, but as the value of the objective function goes down in every step, a return to a previously optimal valuation is not possible.
From a theoretical perspective, it would in particular be interesting to establish non-trivial upper or lower bounds for various pivoting rules. Without such a study, a trivial bound for the proposed approach is provided by the number of strategies (exponential).
A second question is whether this method as whole can be turned into an inner point method [Kar84]. If so, this could be a first step towards showing tractability of discounted payoff games – which would immediately extend to mean-payoff and parity games.
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101032464. It was supported by the EPSRC through the projects EP/X017796/1 (Below the Branches of Universal Trees) and EP/X03688X/1 (TRUSTED: SecuriTy SummaRies for SecUre SofTwarE Development). We also thank Peter Austin from the University of Liverpool for providing an integration of the LP solver.